# Rigged Dynamic Mode Decomposition: Data-Driven Generalized Eigenfunction Decompositions for Koopman Operators

### v1

@@ 4.3 Convergence to the resolvent
As we saw in Example 3.1, the convergence in Eq. 17 need not hold for methods such as EDMD. In particular, convergence in the strong operator topology (the convergence of EDMD) need not imply convergence of the resolvent. The fact that mpEDMD provides a measure-preserving approximation is crucial to Theorem 4.1.
@@ 5 Smooth wave-packet approximations to generalized eigenfunctions
To recover generalized eigenfunctions of $\mathcal{K}$ from sampling the resolvent $\smash{(\mathcal{K}-zI)^{-1}}$, we link $\smash{(\mathcal{K}-zI)^{-1}}$ to generalized eigenfunction in two steps. First, generalized eigenfunctions diagonalize the spectral measure on the rigged Hilbert space. For each Borel subset $S\subset{[{-}\pi,\pi]_{\mathrm{per}}}$ [36],
(18) $\mathcal{E}(S)g=\sum_{k\in\mathcal{I}}\int_{S}\langle g_{\varphi,k}^{*}|g\rangle g_{\varphi,k}{\,\mathrm{d}}\xi_{g_{k}}(\varphi)\qquad\text{for all}\qquad g\in\mathcal{S}.$
@@
The second equality follows from the functional calculus of $\mathcal{K}$. Writing $z=re^{i\theta}$ in modulus-argument form and assuming that $r=(1+\epsilon)>1$, one finds that (c.f. Eqs. 11, 12 and 13)
(20) $\frac{1}{4\pi}\left[F_{\mathcal{E}}(r^{-1}e^{i\theta})-F_{\mathcal{E}}(re^{i\theta})\right]=\frac{1}{2\pi}\int_{{[{-}\pi,\pi]_{\mathrm{per}}}}\frac{(r^{2}-1){\,\mathrm{d}}\mathcal{E}(\varphi)}{1+r^{2}-2r\cos(\theta-\varphi)}=[K^{\rm(P)}_{\epsilon}*\mathcal{E}](\theta).$
@@
Stone’s classic formula [76] recovers $\mathcal{E}$ by taking $r=(1+\epsilon)\downarrow 1$. Similarly, we can combine Eq. 18 and Eq. 20 to recover the generalized Koopman modes and eigenfunctions from $F_{\mathcal{E}}(z)$. We obtain convergence rates for wave-packet approximations to Koopman modes in the continuous spectrum under mild regularity conditions on spectral measures of $\mathcal{K}$. We denote the interior of a closed interval $\mathcal{I}$ by ${\rm int}(\mathcal{I})$ and write $g\in\mathcal{C}^{n,\alpha}(\mathcal{I})$ for a function with $n\in\mathbb{N}\cup\{0\}$ continuous derivatives on $\mathcal{I}$ with an $\alpha$-Hölder continuous $n$th derivative.
@@
In practice, the sublinear convergence rates associated with the Poisson kernel present a computational bottleneck. As noted in [28, Figure 10], as $\epsilon\downarrow 0$, the computational cost typically increases because a larger dictionary of observables is required to compute the resolvent $(\mathcal{K}-zI)^{-1}$ and, consequently, a larger collection of snapshot data is required. Therefore, keeping $\epsilon$ from becoming too small is crucial.
@@ 5.2 Convergence theorems
(28) $\begin{split}g_{\theta}^{\epsilon}=\left[K_{\epsilon}^{\mathbb{T}}\ast\mathcal{E}\right](\theta)g&=\sum_{k\in\mathcal{I}}\int_{[{-}\pi,\pi]_{\mathrm{per}}}\underbrace{K_{\epsilon}^{\mathbb{T}}(\theta-\varphi)}_{\text{sharp peak at $\theta$}}\,\,\underbrace{{\langle g_{\varphi,k}^{*}|g\rangle}}_{\text{K. mode}}\,\,\underbrace{g_{\varphi,k}}_{\text{gen. efun.}}{\,\mathrm{d}}\xi_{g_{k}}(\varphi)\quad\forall g\in\mathcal{S}.\end{split}$
The observable on the left-hand side is a wave-packet approximation of a generalized eigenfunction corresponding to the spectral parameter $\lambda=\exp(i\theta)$, as outlined in Eq. 11. Unlike Poisson wave packets, these wave packets are capable of exploiting smoothness in the spectral measures of $\mathcal{K}$ to provide high-order accurate approximations to the Koopman modes and generalized eigenfunctions associated with the continuous spectrum of $\mathcal{K}$.
@@ 7 Examples
  • Integrable Hamiltonian systems. The associated Koopman operators decompose with a Kronecker structure. The spectrum is continuous on $\mathbb{T}\backslash\{1\}$. Again, we provide an explicit generalized eigenfunction expansion, where the generalized eigenfunctions are plane waves with singular support along hyperplanes of constant energy. As an example, we treat the nonlinear pendulum.
@@ 7.1 Systems with Lebesgue spectrum
The dynamical system $(\Omega,\omega,\mathbf{F})$ with $\omega(\Omega)=1$ has Lebesgue spectrum if there exists an orthonormal basis $\{1,f_{i,j}:i\in\mathcal{I},j\in\mathbb{Z}\}$ of $L^{2}(\Omega,\omega)$ such that
$\mathcal{K}f_{i,j}=f_{i,j-1},\quad i\in\mathcal{I},j\in\mathbb{Z}.$
Here, $\mathcal{I}$ is an index set whose cardinality is called the multiplicity of the Lebesgue spectrum.
@@ 7.1.1 Generalized eigenfunctions of the bilateral shift operator
We can give an explicit rigged Hilbert space for the bilateral shift and, hence, for systems with Lebesgue spectrum. Consider $\mathcal{H}=l^{2}(\mathbb{Z})$ and the bilateral shift operator $\mathcal{L}$ defined by its action on the canonical basis vectors $\mathcal{L}e_{j}=e_{j-1}$, for each $j\in\mathbb{Z}$. As our nuclear space, we take the space of rapidly decreasing sequences [41, p. 85]
@@ 7.2 Integrable Hamiltonian systems
Using the relation (30), we see that $u_{\theta}^{(\mathbf{k},m,j)}$ is a generalized eigenfunction corresponding to $\lambda=e^{i\theta}$. Formally, we have
@@ 7.3 The Lorenz system
where $c$ is a constant chosen so that $g$ is orthogonal to the one-dimensional space spanned by constant functions. The value of $c\approx-0.5382$ is computed by taking the average of $\tanh((xy-5z)/10)$ over our trajectory. Shifting by $c$ ensures that the spectral measure $\xi_{g}$ is purely continuous. Fig. 2 shows the spectral measure computed using $N=1000$, the $6$th order kernel in Fig. 4 and a smoothing parameter $\epsilon=0.05$. There are four noticeable peaks at $\theta\approx 0.414,0.497,0.828,$ and $0.994$. Since the spectral measure is continuous, these do not correspond to eigenvalues. Instead, we can compute generalized eigenfunctions corresponding to these values to obtain approximately coherent structures. These are also shown in Fig. 2. The function labeled (a) is similar to the local spectral projections in [53, Figure 13] (see also [24, Figure 4]), which the authors attributed to an almost-periodic motion of the $z$ component during the time that the state resides in either of the two lobes of the Lorenz attractor. The function labeled (c) corresponds to double the angle $\theta$ of that of (a) and roughly the square of the corresponding function. The structures in (b) and (d) are different and include oscillations within each lobe.
@@ 7.4 Lid-driven cavity flow
Fig. 9 shows the spectral measures of the total kinetic energy, computed using Algorithm 1 with a 10th order kernel, $\epsilon=0.001$ and a dictionary formed by time delays. At $Re=13000$, we see pure point spectrum built from a base frequency. As the Reynolds number increases, the spectrum becomes mixed and then continuous. Figs. 10, 11, 12 and 13 show the generalized Koopman modes computed using Algorithm 2 with a 6th order kernel and $\epsilon=0.01$ and a dictionary of $10$, $100$, $200$ and $500$ POD modes for $Re=13000$, $16000$, $19000$ and $30000$, respectively. As our values of $\theta$, we pick the Strouhal numbers corresponding to the four peaks visible in the spectral measure at $Re=13000$. We have shown the results with the noisy data and the noise-free data. As expected, larger frequencies are more affected by noise. Nevertheless, we are able to compute these modes over the continuous spectrum in the presence of large noise. There are three reasons for robustness of Rigged DMD. The first is that mpEDMD is equivalent to a constrained total least squares problem (see the discussion in [23, Section 6.2]). The second is the smoothing parameter (high-order kernels help, see the discussion in [28, Section 5.1.2]). The third is the averaging procedure in Algorithm 2.
@@ 8 Conclusions
We have developed an algorithm that computes generalized eigenfunction expansions (mode decompositions) of Koopman operators, even in the presence of continuous spectra. Named Rigged DMD, the algorithm leverages Gelfand’s theorem, which asserts that a complete set of generalized eigenfunctions exists within a rigged Hilbert space. Rigged DMD employs unitary approximations of Koopman operators, computed using mpEDMD, to determine the Koopman resolvent. This resolvent is then used to compute smoothed approximations of generalized eigenfunctions for any spectral parameter. We have not only shown examples of rigged Hilbert spaces but also demonstrated how to realize this assumption for general unitary Koopman operators using appropriately weighted vectors derived from delay embedding. In addition to proving convergence theorems for generalized eigenfunctions and spectral measures, we have demonstrated the algorithm’s applicability and robustness through several examples. There is considerable potential for future work, and we briefly discuss two avenues. The first involves applying Rigged DMD to reduced order models and control problems involving continuous spectra. Specifically, Rigged DMD opens the door to control over the interactions between modes corresponding to the continuous and discrete spectra. Second, we relied on the Koopman operator being unitary (i.e., the system is measure-preserving and invertible). We aim to extend Rigged DMD to non-unitary and non-normal Koopman operators, a development that would require adapting these techniques to scenarios where the nuclear spectral theorem does not apply.
@@ Appendix B Construction of high-order periodic kernels
where $\rho(x)=e^{inx}-1$. Since $\rho$ is a bounded, smooth function, and $\rho(0)=0$, we may apply [19, Theorem 5.2] to deduce that Eq. 21 holds.

### v2

@@ 4.3 Convergence to the resolvent
As we saw in Example 3.1, the convergence in Eq. 17 need not hold for methods such as EDMD. In particular, convergence in the strong operator topology (the convergence of EDMD) need not imply convergence of the resolvent. The fact that mpEDMD provides a unitary approximation is crucial to Theorem 4.1.
There are many natural settings where the dictionary does not span all of $L^{2}(\Omega,\omega)$ (for example, see Section 7.1.2). Rather, in this setting the observable of interest belongs to a closed $\mathcal{K}$-invariant subspace of $L^{2}(\Omega,\omega)$ and the dictionary spans this space in an appropriate limit. Theorem 4.1 still holds if $L^{2}(\Omega,\omega)$ is replaced by any closed $\mathcal{K}$-invariant subspace thereof in both the hypotheses and the conclusions. That is, if $V_{N}$ spans this subspace in the limit $N\rightarrow\infty$, then the mpEDMD resolvent must also converge to the resolvent restricted to that subspace in the strong operator topology. Consequently, we can use the mpEDMD resolvent to approximate generalized Koopman modes for any observable in a closed $\mathcal{K}$-invariant subspace.
@@ 5 Smooth wave-packet approximations to generalized eigenfunctions
To recover generalized eigenfunctions of $\mathcal{K}$ by sampling the resolvent $\smash{(\mathcal{K}-zI)^{-1}}$, we link $\smash{(\mathcal{K}-zI)^{-1}}$ to generalized eigenfunction in two steps. First, generalized eigenfunctions diagonalize the spectral measure on the rigged Hilbert space. For each Borel subset $S\subset{[{-}\pi,\pi]_{\mathrm{per}}}$ [45],
(18) $\mathcal{E}(S)g=\sum_{k\in\mathcal{J}}\int_{S}\langle g_{\varphi,k}^{*}|g\rangle g_{\varphi,k}{\,\mathrm{d}}\xi_{g_{k}}(\varphi)\qquad\text{for all}\qquad g\in\mathcal{S}.$
@@
The second equality follows from the functional calculus of $\mathcal{K}$ [^4] Carathéodory functions play a key role in the theory of orthogonal polynomials on the unit circle [85], where they are the CMV analogue of the so called $m$-function ($z\mapsto\langle(J-z)^{-1}e_{1},e_{1}\rangle$ for a matrix $J$) in the theory of Jacobi matrices. For connections between the two, see [86, Section 2.3]. Writing $z=re^{i\theta}$ in modulus-argument form and assuming that $r=1/(1+\epsilon)<1$, one finds that (c.f. Eqs. 11, 12 and 13)
(20) $\frac{1}{4\pi}\left[F_{\mathcal{E}}(re^{i\theta})-F_{\mathcal{E}}(r^{-1}e^{i\theta})\right]=\frac{1}{2\pi}\int_{{[{-}\pi,\pi]_{\mathrm{per}}}}\frac{(1-r^{2}){\,\mathrm{d}}\mathcal{E}(\varphi)}{1+r^{2}-2r\cos(\theta-\varphi)}=[K^{\rm(P)}_{\epsilon}*\mathcal{E}](\theta).$
@@
A measure on ${[{-}\pi,\pi]_{\mathrm{per}}}$ is the weak limit of the real part of its Carathéodory function as $r\uparrow 1$ and Stone’s classic formula [88] recovers $\mathcal{E}$ by taking $r\uparrow 1$. Similarly, we can combine Eq. 18 and Eq. 20 to recover the generalized Koopman modes and eigenfunctions from $F_{\mathcal{E}}(z)$. We obtain convergence rates for wave-packet approximations to Koopman modes in the continuous spectrum under mild regularity conditions on spectral measures of $\mathcal{K}$. We denote the interior of a closed interval $\mathcal{I}$ by ${\rm int}(\mathcal{I})$ and write $g\in\mathcal{C}^{n,\alpha}(\mathcal{I})$ for a function with $n\in\mathbb{N}\cup\{0\}$ continuous derivatives on $\mathcal{I}$ with an $\alpha$-Hölder continuous $n$th derivative.
@@
In practice, the sublinear convergence rates associated with the Poisson kernel present a computational bottleneck. As noted in [29, Figure 10], as $\epsilon\downarrow 0$, the computational cost typically increases because a larger dictionary of observables is required to compute the resolvent $(\mathcal{K}-zI)^{-1}$ and, consequently, a larger collection of snapshot data is required. Therefore, keeping $\epsilon$ from becoming too small is crucial.
@@ 5.2 Convergence theorems
(28) $\begin{split}g_{\theta}^{\epsilon}=\left[K_{\epsilon}^{\mathbb{T}}\ast\mathcal{E}\right](\theta)g&=\sum_{k\in\mathcal{J}}\int_{[{-}\pi,\pi]_{\mathrm{per}}}\underbrace{K_{\epsilon}^{\mathbb{T}}(\theta-\varphi)}_{\text{sharp peak at $\theta$}}\,\,\underbrace{{\langle g_{\varphi,k}^{*}|g\rangle}}_{\text{K. mode}}\,\,\underbrace{g_{\varphi,k}}_{\text{gen. efun.}}{\,\mathrm{d}}\xi_{g_{k}}(\varphi)\quad\forall g\in\mathcal{S}.\end{split}$
The observable on the left-hand side is a wave-packet approximation of a generalized eigenfunction corresponding to the spectral parameter $\lambda=\exp(i\theta)$, as outlined in Eq. 11. Unlike Poisson wave packets, these wave packets are capable of exploiting smoothness in the spectral measures of $\mathcal{K}$ to provide high-order accurate approximations to the Koopman modes and generalized eigenfunctions associated with the continuous spectrum of $\mathcal{K}$. Beyond improved convergence, there are as least two other advantages of high-order kernels [29, Section 5.1]: an improved stability to noise with suitably selected $\epsilon$; and improved localization of the kernel.
@@ 7 Examples
  • Integrable Hamiltonian systems. The associated Koopman operators decompose with a Kronecker structure. The spectrum is continuous on $\mathbb{T}\backslash\{1\}$. Again, we provide an explicit generalized eigenfunction expansion, where the generalized eigenfunctions are plane waves with singular support along hyperplanes of constant energy (with connections to the Radon and Fourier transforms for the action and angle parts, respectively). As an example, we treat the nonlinear pendulum.
@@ 7.1 Systems with Lebesgue spectrum
The dynamical system $(\Omega,\omega,\mathbf{F})$ with $\omega(\Omega)=1$ has Lebesgue spectrum if there exists an orthonormal basis $\{1,f_{i,j}:i\in\mathcal{J},j\in\mathbb{Z}\}$ of $L^{2}(\Omega,\omega)$ such that
$\mathcal{K}f_{i,j}=f_{i,j-1},\quad i\in\mathcal{J},j\in\mathbb{Z}.$
Here, $\mathcal{J}$ is an index set whose cardinality is called the multiplicity of the Lebesgue spectrum.
@@ 7.1.1 Generalized eigenfunctions of the bilateral shift operator
We can give an explicit rigged Hilbert space for the bilateral shift and, hence, for systems with Lebesgue spectrum. Consider $\mathcal{H}=l^{2}(\mathbb{Z})$ and the bilateral shift operator $\mathcal{L}$ defined by its action on the canonical basis vectors $\mathcal{L}e_{j}=e_{j-1}$, for each $j\in\mathbb{Z}$. As our nuclear space, we take the space of rapidly decreasing sequences [50, p. 85]
@@ 7.2 Integrable Hamiltonian systems
This integral is the Radon transform of $\hat{g}_{\mathbf{k}}(\mathbf{I})$ [40, Chapter 6] with an additional integration against the function $\overline{\psi_{j}^{(\mathbf{k})}(Q_{\mathbf{k}}(\mathbf{I}))}$. Using the relation (30), we see that $u_{\theta}^{(\mathbf{k},m,j)}$ is a generalized eigenfunction corresponding to $\lambda=e^{i\theta}$. Formally, we have
@@ 7.3 The Lorenz system
where $c$ is a constant chosen so that $g$ is orthogonal to the one-dimensional space spanned by constant functions. The value of $c\approx-0.5382$ is computed by taking the average of $\tanh((xy-5z)/10)$ over our trajectory. Shifting by $c$ ensures that the spectral measure $\xi_{g}$ is purely continuous. Fig. 2 shows the spectral measure computed using $N=1000$, the $6$th order kernel in Fig. 4 and a smoothing parameter $\epsilon=0.05$. There are four noticeable peaks at $\theta\approx 0.414,0.497,0.828,$ and $0.994$. Since the spectral measure is continuous, these do not correspond to eigenvalues. Instead, we can compute generalized eigenfunctions corresponding to these values to obtain approximately coherent structures. These are also shown in Fig. 2. The function labeled (a) is similar to the local spectral projections in [63, Figure 13] (see also [25, Figure 4]), which the authors attributed to an almost-periodic motion of the $z$ component during the time that the state resides in either of the two lobes of the Lorenz attractor. The function labeled (c) corresponds to double the angle $\theta$ of that of (a) and roughly the square of the corresponding function. The structures in (b) and (d) are different and include oscillations within each lobe.
@@ 7.4 Lid-driven cavity flow
Fig. 9 shows the spectral measures of the total kinetic energy, computed using Algorithm 1 with a 10th order kernel, $\epsilon=0.001$ and a dictionary formed by time delays. At $Re=13000$, we see pure point spectrum built from a base frequency. As the Reynolds number increases, the spectrum becomes mixed and then continuous. Figs. 10, 11, 12 and 13 show the generalized Koopman modes computed using Algorithm 2 with a 6th order kernel and $\epsilon=0.01$ and a dictionary of $10$, $100$, $200$ and $500$ POD modes for $Re=13000$, $16000$, $19000$ and $30000$, respectively. As our values of $\theta$, we pick the Strouhal numbers corresponding to the four peaks visible in the spectral measure at $Re=13000$. We have shown the results with the noisy data and the noise-free data. As expected, larger frequencies are more affected by noise. Nevertheless, we are able to compute these modes over the continuous spectrum in the presence of large noise. There are three reasons for robustness of Rigged DMD. The first is that mpEDMD is equivalent to a constrained total least squares problem (see the discussion in [24, Section 6.2]). The second is the smoothing parameter (high-order kernels help, see the discussion in [29, Section 5.1.2]). The third is the averaging procedure in Algorithm 2.
@@ 8 Conclusions
We have developed an algorithm that computes generalized eigenfunction expansions (mode decompositions) of Koopman operators, even in the presence of continuous spectra. Named Rigged DMD, the algorithm leverages Gelfand’s theorem, which asserts that a complete set of generalized eigenfunctions exists within a rigged Hilbert space. Rigged DMD employs unitary approximations of Koopman operators, computed using mpEDMD, to determine the Koopman resolvent. This resolvent is then used to compute smoothed approximations of generalized eigenfunctions for any spectral parameter. We have not only shown examples of rigged Hilbert spaces but also demonstrated how to realize this assumption for general unitary Koopman operators using appropriately weighted vectors derived from delay embedding. In addition to proving convergence theorems for generalized eigenfunctions and spectral measures, we have demonstrated the algorithm’s applicability and robustness through several examples.
There is considerable potential for future work, and we briefly discuss two avenues. The first involves applying Rigged DMD to reduced order models and control problems involving continuous spectra. Specifically, Rigged DMD opens the door to control over the interactions between modes corresponding to the continuous and discrete spectra. Second, we relied on the Koopman operator being unitary (i.e., the system is measure-preserving and invertible). We aim to extend Rigged DMD to non-unitary and non-normal Koopman operators, a development that would require adapting these techniques to scenarios where the nuclear spectral theorem does not apply.
Finally, several methods now exist for computing spectral measures of Koopman operators. It would be interesting to explore which of these methods can be extended to compute generalized eigenfunctions. Furthermore, we believe the community is at a stage where comparing these approaches for spectral measures would be beneficial, helping to establish a guideline for identifying the most suitable methods for different types of problems.
@@ Appendix B Construction of high-order periodic kernels
where $\rho(x)=e^{inx}-1$. Since $\rho$ is a bounded, smooth function, and $\rho(0)=0$, we may apply [20, Theorem 5.2] to deduce that Eq. 21 holds.